Hank deposited $750 in the bank for 3 years. He earned $135 in simple interest at the end of the three years. What was the annual interest rate?
Answer by rfer(16322)

I=Prt
r=I/Pt
r=135/(750*3)
r=135/2250
r=0.06=6%